OpenStudy (anonymous):

Integrate this function on the domain from [0,(Pi/2)] where f(x) = 2SinxCosx[3Sinx-4Sinx^2(x)]dx

OpenStudy (freckles):

you should try to let u=sin(x)

OpenStudy (freckles):

\[\int\limits_{\sin(0)}^{\sin(\frac{\pi}{2})} 2 u (3u-4u^2) du\]

OpenStudy (freckles):

multiply then integrate then plug in limits